OpenStudy (anonymous):

Is 1/sin^4 x the same as csc^4 x?

OpenStudy (kinggeorge):

Yes, since \[{1\over{\sin^4 (x)}} = \left({1\over \sin(x)}\right)^4\]and \(\csc(x) = {1\over \sin(x)}\)

OpenStudy (anonymous):

thanks